OpenStudy (anonymous):

FAN AND MEDAL!!!! Mickey bought 3 books from the bargain bin at his school's book fair and paid nine dollars. If buying books from the bargain bin follows a direct variation, how much did his sister pay for the 4 books she bought? seven dollars ten dollars twelve dollars sixteen dollars

OpenStudy (anonymous):

) A direct variation means that the books do not change price. Mickey bought 3 book for $9 total, the books being $3 each. If his sister buys 4 books, she will pay 4 x 3 = $12, C 2) There are 18 friends. If 4/9 of them chose song A, we have to divide 18 by the denominator, 9, and then multiply that answer by the numerator, 4. 18/9 = 2. 2 x 4 = 8. 8 friends chose song A. This leaves 10 friends who have chosen other songs. If 1/5 of those 10 friends chose song B, we use the same process to find out how many friends chose song B. 10/5 = 2. 2 x 1 = 2. 2 friends chose song B. This means that (8 + 2 = 10) 10 friends chose songs A and B. 18 - 10 shows that there are 8 friends that didn't choose song A or B, which means that they chose song C. A 3) Just look at the math on Gustav's steps. His first step checks out completely. 72 - 32 does = 40. However, his second step doesn't look so promising. If he uses front-end estimation in every step, 0.93 would round down to 0, and so would 0.07, making 0.93 - 0.07 = 0, rather than 1. B

OpenStudy (anonymous):

hope i helped

OpenStudy (anonymous):

now wheres my medal and fan that u said u would give

OpenStudy (anonymous):

you type fast

OpenStudy (anonymous):

thx

OpenStudy (uri):

@Matt0918 Do you really think he typed this in 3 seconds? LMAOOOOOOOOOO

OpenStudy (anonymous):

its called yahoo

OpenStudy (anonymous):

look it up

OpenStudy (anonymous):

ok I thought he typed it sorry

OpenStudy (anonymous):

lmao